OpenStudy (anonymous):

What are the amplitude, period, phase shift, and midline of f(x) = 7 cos(2x + π) − 3?

OpenStudy (anonymous):

Amplitude = −3; period: π; phase shift: x = negative pi over 2; midline: y = 3 Amplitude: 7; period: π; phase shift: x = negative pi over 2; midline: y = −3 Amplitude: 7; period: 2π; phase shift: x = pi over 2; midline: y = 3 Amplitude: −3; period: 2π; phase shift: x = pi over 2; midline: y = −3

OpenStudy (aum):

If f(x) = Acos(Bx + C) + D, then: Amplitude = A Period = (2*pi) / B Phase shift = -C / B Midline is: y = D

OpenStudy (anonymous):

Okay, so would I use that formula for all of them?

OpenStudy (aum):

Yes. And the same formula applies if it is sine function instead of cosine.

OpenStudy (anonymous):

okay I'll see what I get then let you know

OpenStudy (aum):

Alright.

OpenStudy (anonymous):

how do i find x?

OpenStudy (aum):

Not sure what you mean? For this problem you don't have to solve for x.

OpenStudy (anonymous):

But like in the formula thing there's x...

OpenStudy (anonymous):

You probably think i'm really stupid ehhh sorry :(

OpenStudy (aum):

You just compare the given function to the general function and identify the constants A, B, C and D and then plug it into the formula: f(x) = 7 cos(2x + π) − 3 Compare this to the general form, f(x) = Acos(Bx + C) + D A = 7 B = 2 C = π D = -3 Amplitude = A = 7 Period = 2π / B = 2π / 2 = π ....

OpenStudy (anonymous):

OHHH okay thank you so muchhhh sorry for burdening you with all of my math problems I seriously do appreciate it :)

OpenStudy (anonymous):

C is the correct answer right?

OpenStudy (aum):

Hey, no problem at all. We are here to assist. Not C. C has a period of 2π. We calculated the period above to be π. Amplitude = A = 7 Period = 2π / B = 2π / 2 = π To finish the other two: Phase shift = -C / B = -π / 2 Midline is: y = D; y = -3

OpenStudy (anonymous):

Okay. It's A?

OpenStudy (anonymous):

?!

OpenStudy (aum):

The answer we got above is: Amplitude: 7; period: π; phase shift: x = -π / 2; midline: y = −3

OpenStudy (aum):

Choice B.

OpenStudy (anonymous):

I LOVE YOU

OpenStudy (aum):

lol